OpenStudy (anonymous):

Use integration by parts to evaluate the integral. 343x^2cos(7x)dx

OpenStudy (anonymous):

u = 343x^2 dv = cos(7x)dx du = 686xdx v = (sin(7x))/7 ∫udv = uv - ∫vdu = 343x^2*(sin(7x))/7 - ∫(sin(7x))/7*686xdx =49x^2*sin(7x) - ∫sin(7x)*98xdx Then do parts again on the second term, choosing the power function for u...

OpenStudy (turingtest):

49x^2(sin7x)+14cos7x-2sin7x

OpenStudy (turingtest):

i think

OpenStudy (turingtest):

hooray I was right!